# Towards Trustworthy Reranking: A Simple yet Effective Abstention Mechanism

### v4

@@ 5.6 Computational Overhead Analysis
Figure 5: Computational overhead analysis. Comparison between relevance score ($g_{\bm{\theta}}$) and confidence score ($u_{\text{lin}}$) calculation times (on average, 8 ms and 0.1 ms respectively on one Apple M1 CPU) (model: all-MiniLM-L6-v2, dataset: StackOverflow).
@@
Our data-based abstention method is free of any significant time overhead. $u_{\text{lin}}$-based confidence estimation accounts for only 1.2% of the time needed for relevance scores computation (Figure 5), making it an almost cost-free option [^10] If we used a cross-encoder, documents could not be pre-embedded as they need to be appended to the input query. In that scenario, the computation time required for $g_{\bm{\theta}}$ would therefore increase, while that required for $u_{\text{lin}}$ would remain unchanged, thus reducing the relative extra running time..
@@ 6 Conclusion
We developed a lightweight abstention mechanism tailored to real-world constraints, agnostic to the ranking method and shown to be effective across various settings and use cases. Our method, which only requires black-box access to relevance scores and a small reference set for calibration, is essentially a free-lunch to plug to any retrieval system, in order to gain better control over the whole IR pipeline. A logical next step would be to generalize findings to ensembling-based retrieval methods which boost downstream performance while outputting a richer signal for confidence estimations.
Limitations
Our work studies abstention mechanisms within the scope of the reranking stage of a larger information retrieval pipeline. While this enables estimating confidence for the final prediction of the IR system, the lack of existing document-wise-labeled datasets [^11] Datasets that explicitly label relevant and irrelevant documents, which is not the case with classical retrieval benchmarks (e.g., BEIR Thakur et al. (2021)). has prevented us from evaluating our methods at the primary ranking stage, although they should theoretically be directly applicable. Additionally, our study primarily focuses on reranking scenarios with the top-10 retrieved documents, which correspond to standard industrial use cases, but it would be interesting to extend this scenario to top-50 or top-100 documents. Our linear-regression-based abstention method is easily scalable and should theoretically benefit from additional document scores in its decision-making process, but the performance impact remains to be quantified. Finally, in RAG applications, an LLM is given the retrieved documents as context, but might not use all of them to formulate an answer, thus acting as a final abstention mechanism. Studies have even shown these RAG LLMs work best when given a certain ratio of non-relevant documents in order to induce a filtering behavior Cuconasu et al. (2024). Studying the impact of abstention mechanisms on end-to-end RAG performance would thus be an interesting future work direction. Additionally, in the future, we could rely on anomaly detection algorithms to remove any need for training, specifically, we want to investigate the use of isolation forest Darrin et al. (2023), data depths Colombo et al. (2022, 2023b); Staerman et al. (2021); Picot et al. (2022a, b), information measures Colombo et al. (2023a) and projections Darrin et al. (2022); Picot et al. (2022c), and Wasserstein distances Guerreiro et al. (2022).
@@ Acknowledgements
Training compute is obtained on the Jean Zay supercomputer operated by GENCI IDRIS through compute grant 2023-AD011014668R1, AD010614770 as well as on Adastra through project c1615122, cad15031, cad14770.
Appendices
@@ Appendix A Additional Information on Models and Datasets
Table 6: Dataset details after preprocessing.
$N$ $k$ $p_{\text{min}}$ $p_{\text{max}}$ $\bar{p}$
SciDocs 3978 10 1 5 4.9
AskUbuntu 351 10 1 5 3.6
StackOverflow 2980 10 1 5 1.2
Alloprof 2316 10 1 5 1.3
CMedQAv1 1000 10 1 5 1.8
Mmarco 100 10 1 3 1.1
Table 7: Details on raw datasets.
Reference Language $N$ $k_{\text{min}}$ $k_{\text{max}}$ $\bar{k}$ $p_{\text{min}}$ $p_{\text{max}}$ $\bar{p}$
SciDocs Cohan et al. 2020 English 3978 26 60 29.8 1 10 4.9
AskUbuntu Lei et al. 2015 English 375 20 20 20.0 1 20 6.0
StackOverflow Zhang et al. 2015 English 2992 20 30 29.9 1 6 1.2
Alloprof Lefebvre-Brossard et al. 2023 French 2316 10 35 10.7 1 29 1.3
CMedQAv1 Zhang et al. 2017 Chinese 1000 100 100 100.0 1 19 1.9
Mmarco Bonifacio et al. 2021 Chinese 100 1000 1002 1000.3 1 3 1.1
@@
msmarco-MiniLM-L6-cos-v5 Reimers and Gurevych 2021 English bi cos 23
msmarco-distilbert-dot-v5 Reimers and Gurevych 2021 English bi cos 66
ms-marco-TinyBERT-L-2-v2 Reimers and Gurevych 2021 English cross n.a. 4
ms-marco-MiniLM-L-6-v2 Reimers and Gurevych 2021 English cross n.a. 23
stsb-TinyBERT-L-4 Reimers and Gurevych 2019 English cross n.a. 14
stsb-distilroberta-base Reimers and Gurevych 2019 English cross n.a. 82
multi-qa-distilbert-cos-v1 Reimers and Gurevych 2019 English bi cos 66
multi-qa-MiniLM-L6-cos-v1 Reimers and Gurevych 2019 English bi cos 23
all-MiniLM-L6-v2 Reimers and Gurevych 2019 English bi cos 23
all-distilroberta-v1 Reimers and Gurevych 2019 English bi cos 82
all-mpnet-base-v2 Reimers and Gurevych 2019 English bi cos 109
quora-distilroberta-base Reimers and Gurevych 2019 English cross n.a. 82
qnli-distilroberta-base Reimers and Gurevych 2019 English cross n.a. 82
@@
In this section, we give details on the datasets used in our work, before the preprocessing phase (Table 7), and after preprocessing (Table 6). As in the main text, $N$ denotes the number of instances and $k$ represents the number of candidate documents. In the same logic, $k_{\text{min}}$, $k_{\text{max}}$ and $\bar{k}$ respectively denote the minimum, maximum and average number of candidate documents in the dataset. In addition, $p_{\text{min}}$, $p_{\text{max}}$, $\bar{p}$ denote the minimum, maximum and average number of positive documents respectively. Additionally, we provide information on models’ raw performance (without abstention) across the whole benchmark in Table 9.
@@ Appendix C Additional Results on Domain Adaptation
Figure 6: Domain adaptation study. Coefficients of the $u_{\text{lin}}$ confidence function (model: ember-v1, metric: mAP). Recall that documents are sorted in increasing order of relevance scores. Therefore, document 10 corresponds to the document with highest relevance score while document 1 is the least relevant to the query.
Table 11: Extensive domain adaptation study. nAUCs (in %) for all reference-test relevant pairs (model: ember-v1, metric: mAP).
@@
A simple observation that can be made from Figure 6 is that the value taken by the coefficients depends very strongly on the number of positive documents per instance (see Table 6). Indeed, when we look at the results for the SciDocs dataset (five positive documents per instance on average), we see that $u_{\text{lin}}$ has positive weights on the first five documents and negative weights on the last five. A similar observation can be made for the other five datasets.
@@ Appendix D Additional Reference-Based Confidence Functions
In this section, we introduce two other ways to tackle confidence assessment in a reference-based setting: classification-based and distance-based methods.
D.1 Classification-Based Approach
@@
$\begin{split}\mathcal{S}_{\bm{\theta}}^{\text{clf}}=\bigr{\{}\left(\mathbf{z},\mathbbm{1}_{\left\{y>m^{+}\right\}}-\mathbbm{1}_{\left\{y<m^{-}\right\}}\right)\ |\ \\ (\mathbf{z},y)\in\mathcal{S}_{\bm{\theta}}\bigl{\}}.\end{split}$ (9)
where $m^{+}>m^{-}\in\mathbb{R}$ are the thresholds above and below which an instance is considered good and bad respectively. Then, we fit a probabilistic classifier $\pi:\mathbb{R}^{k}\rightarrow[0,1]^{3}$ on $\mathcal{S}_{\bm{\theta}}^{\text{clf}}$, taking a vector of sorted scores $s(\mathbf{z})\in\mathbb{R}^{k}$ as input and returning the estimated probabilities of $s(\mathbf{z})$ to belong to class $-1$ (bad instances), 00 (average instances) and $1$ (good instances). We finally define confidence function $u_{\text{clf}}$ such that for a given unsorted vector of relevance scores $\mathbf{z}$, $u_{\text{clf}}(\mathbf{z})=\pi_{2}\left(s(\mathbf{z})\right)-\pi_{0}\left(s(\mathbf{z})\right)$ (probability that $\mathbf{z}$ stems from a good instance minus the probability that $\mathbf{z}$ stems from a bad instance). The higher $u_{\text{clf}}(\mathbf{z})$, the more confident we are that $\mathbf{z}\in\mathbb{R}^{k}$ stems from a good instance.
In this work, we use a logistic regressor Cox and Snell (1989) with $l_{2}$ regularization parameter $\lambda=0.1$, $u_{\text{log}}$, and consider by default that the top and bottom 25% of instances with respect to metric function $m$ represent the good and bad instances respectively. We challenge this parameter in Appendix D.4.
D.2 Distance-Based Approach
@@
For this purpose, based on thresholds $m^{+}$ and $m^{-}$ defined in Section D.1, we build $\mathcal{Z}_{\bm{\theta}}^{+}$ and $\mathcal{Z}_{\bm{\theta}}^{-}$, the sets composed of good and bad instances respectively. Formally,
@@
Then, we construct a confidence scorer $u_{\text{dist}}$ based on the distributions observed in $\mathcal{Z}_{\bm{\theta}}^{+}$ and $\mathcal{Z}_{\bm{\theta}}^{-}$. We define $u_{\text{dist}}=\delta\left(s(\cdot),\mathcal{Z}_{\bm{\theta}}^{-}\right)-\delta\left(s(\cdot),\mathcal{Z}_{\bm{\theta}}^{+}\right)$ as the difference between the distance to the set of bad instances and the distance to the set of good instances [^12] As for the other reference-based $u$ functions, $u_{\text{dist}}$ takes a vector of unsorted scores as input, sorting being encompassed inside the function.. Intuitively, the higher $u_{\text{dist}}(\mathbf{z})$, the more confident we are that $\mathbf{z}$ stems from a good instance.
In this experiment, we rely on the Mahalanobis distance Mahalanobis (2018); Lee et al. (2018); Ren et al. (2021). We compute the Mahalanobis distance between set $\mathcal{Z}\subset\mathbb{R}^{k}$ and $\mathbf{z}\in\mathbb{R}^{k}$ as
@@
where $\bm{\mu}_{\mathcal{Z}}\in\mathbb{R}^{k}$ denotes the average vector of reference set $\mathcal{Z}$ and $\Sigma_{\mathcal{Z}}^{-1}\in\mathbb{R}^{k\times k}$ the inverse of its covariance matrix.
@@ D.3 Experimental Results
Table 12: Abstention performance assessment. nAUCs (in %) averaged over the whole set of available models, for each method including $u_{\text{log}}$ and $u_{\text{mah}}$.
Method $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{log}}$ $u_{\text{mah}}$
@@
SciDocs mAP 49.2 58.7 4.7 65.4 68.8 65.8
mNDCG 54.9 62.9 10.2 66.6 70.2 66.8
mRR 80.2 80.3 41.3 80.5 78.1 65.6
AskUbuntu mAP 24.1 16.2 7.3 22.0 20.0 10.1
mNDCG 28.1 15.7 8.6 24.5 21.9 10.5
mRR 32.7 16.6 14.0 26.2 24.2 11.3
StackOverflow mAP 20.7 24.6 35.7 42.6 42.0 37.5
mNDCG 21.0 25.0 35.6 42.6 42.1 37.4
mRR 21.4 24.8 36.1 42.8 42.3 37.7
Alloprof mAmP 16.6 17.0 19.3 29.7 29.4 25.1
mNDCG 16.8 16.9 18.8 29.3 28.9 24.9
mRR 16.9 16.7 19.1 28.9 28.7 24.3
CMedQAv1 mAP 24.4 23.3 22.3 30.6 29.1 25.7
mNDCG 25.0 23.5 22.3 30.7 27.0 24.8
mRR 26.5 24.2 23.4 31.5 30.3 25.5
Mmarco mAP 30.1 31.0 39.4 34.0 24.9 22.7
mNDCG 30.5 30.9 38.8 34.1 24.9 23.3
mRR 30.1 31.0 39.6 34.3 23.5 22.8
Average mAP 27.5 28.5 21.5 37.4 35.7 31.1
mNDCG 29.4 29.1 22.4 38.0 35.8 31.3
mRR 34.6 32.3 28.9 40.7 37.8 31.2
@@
D.4 Impact of Instance Qualification
Figure 7: Instance qualification analysis. nAUC vs. instance quantile threshold for all datasets (model: ember-v1, metric: mAP).
In this section, we analyze the impact of the instance qualification threshold on the performance of the $u_{\text{log}}$ and $u_{\text{mah}}$ confidence functions. More precisely, the idea of this experiment is to vary the qualification threshold [^13] As a reminder, a threshold of 10% means that good instances represent the top-10% with regard to the metric of interest, while bad instances represent the bottom-10%. and to test the impact on nAUC over the whole benchmark. For the sake of readability, we present the results for the ember-v1 model and for the mAP only (see Figure 7).

### v5

@@ 5.6 Computational Overhead Analysis
Figure 2: Comparison between relevance ($g_{\bm{\theta}}$) and confidence ($u_{\text{lin}}$) assessment times on one Apple M1 CPU (model: all-MiniLM-L6-v2, dataset: StackOverflow).
@@
Our data-based abstention method is free of any significant time overhead. $u_{\text{lin}}$-based confidence estimation accounts for only 1.2% of relevance scores calculation time (Figure 2), making it an almost cost-free option. [^11] If we used a cross-encoder, documents could not be pre-embedded as they need to be appended to the input query. In that scenario, the computation time required for $g_{\bm{\theta}}$ would therefore increase, while that required for $u_{\text{lin}}$ would remain unchanged, thus reducing the relative extra running time.
@@ 6 Conclusion
We developed a lightweight abstention mechanism tailored to real-world constraints, agnostic to the ranking method and shown to be effective across various settings and use cases. Our method, which only requires black-box access to relevance scores and a small reference set for calibration, is essentially a free-lunch to plug to any retrieval system, in order to gain better control over the whole IR pipeline.
Limitations and Future Work
Our study investigates abstention mechanisms specifically within the reranking stage of an IR pipeline. While this provides a framework for estimating the confidence of the final prediction, the absence of datasets with explicit document-wise labels has limited our ability to evaluate these methods in the primary retrieval stage. These datasets typically lack labels indicating relevant and irrelevant documents, as is the case with classical retrieval benchmarks like BEIR (Thakur et al., 2021). However, in theory, our methods should be applicable at this earlier stage of the pipeline, making this a potential area for future exploration.
Moreover, our experiments have mainly focused on reranking the top-10 retrieved documents, aligning with common industrial use cases. A natural extension would be to explore how abstention mechanisms perform when applied to larger sets of documents, such as the top-50 or top-100 results. Since our linear-regression-based abstention method is designed to scale efficiently, it could potentially benefit from the richer decision-making context provided by more document scores, although this impact has yet to be quantified.
A broader consideration, however, lies in the potential of generative models, such as those used in RAG, to handle abstention on their own (Zhang et al., 2024). These models might inherently possess the capability to estimate confidence without needing a separate abstention mechanism, a compelling prospect for further investigation.
Another promising avenue is to test abstention mechanisms on relevance-scoring methods that offer more detailed information, such as token-level relevance scores for each query-document pair. Methods like ColBERT (Khattab & Zaharia, 2020) or the ColPALI (Faysse et al., 2024) model family provide richer signals, which could lead to improved abstention performance. It would be interesting to see if the additional granularity in relevance scores results in better decision-making, particularly in scenarios where token-level understanding could refine the abstention process.
@@ Acknowledgements
Training compute was obtained on the Jean Zay supercomputer operated by GENCI IDRIS through compute grant 2023-AD011014668R1, AD010614770 as well as on Adastra through project c1615122, cad15031, cad14770.
@@ Appendix A Additional Information on Models and Datasets
[…]
tableDetails on raw datasets. Reference Language $N$ $k_{\text{min}}$ $k_{\text{max}}$ $\bar{k}$ $p_{\text{min}}$ $p_{\text{max}}$ $\bar{p}$ SciDocs Cohan et al. 2020 English 3978 26 60 29.8 1 10 4.9 AskUbuntu Lei et al. 2015 English 375 20 20 20.0 1 20 6.0 StackOverflow Zhang et al. 2015 English 2992 20 30 29.9 1 6 1.2 Alloprof Lefebvre-Brossard et al. 2023 French 2316 10 35 10.7 1 29 1.3 CMedQAv1 Zhang et al. 2017 Chinese 1000 100 100 100.0 1 19 1.9 Mmarco Bonifacio et al. 2021 Chinese 100 1000 1002 1000.3 1 3 1.1
[…]
tableDataset details after preprocessing. $N$ $k$ $p_{\text{min}}$ $p_{\text{max}}$ $\bar{p}$ SciDocs 3978 10 1 5 4.9 AskUbuntu 351 10 1 5 3.6 StackOverflow 2980 10 1 5 1.2 Alloprof 2316 10 1 5 1.3 CMedQAv1 1000 10 1 5 1.8 Mmarco 100 10 1 3 1.1
@@
msmarco-MiniLM-L6-cos-v5 Reimers & Gurevych 2021 English bi cos 23
msmarco-distilbert-dot-v5 Reimers & Gurevych 2021 English bi cos 66
ms-marco-TinyBERT-L-2-v2 Reimers & Gurevych 2021 English cross n.a. 4
ms-marco-MiniLM-L-6-v2 Reimers & Gurevych 2021 English cross n.a. 23
stsb-TinyBERT-L-4 Reimers & Gurevych 2019 English cross n.a. 14
stsb-distilroberta-base Reimers & Gurevych 2019 English cross n.a. 82
multi-qa-distilbert-cos-v1 Reimers & Gurevych 2019 English bi cos 66
multi-qa-MiniLM-L6-cos-v1 Reimers & Gurevych 2019 English bi cos 23
all-MiniLM-L6-v2 Reimers & Gurevych 2019 English bi cos 23
all-distilroberta-v1 Reimers & Gurevych 2019 English bi cos 82
all-mpnet-base-v2 Reimers & Gurevych 2019 English bi cos 109
quora-distilroberta-base Reimers & Gurevych 2019 English cross n.a. 82
qnli-distilroberta-base Reimers & Gurevych 2019 English cross n.a. 82
@@
In this section, we give details on the datasets used in our work, before the preprocessing phase (Table A), and after preprocessing (Table A). As in the main text, $N$ denotes the number of instances and $k$ represents the number of candidate documents. In the same logic, $k_{\text{min}}$, $k_{\text{max}}$ and $\bar{k}$ respectively denote the minimum, maximum and average number of candidate documents in the dataset. In addition, $p_{\text{min}}$, $p_{\text{max}}$, $\bar{p}$ denote the minimum, maximum and average number of positive documents respectively. Additionally, we provide information on models’ raw performance (without abstention) across the whole benchmark in Table 2.
@@ Appendix C Additional Results on Domain Adaptation
Figure 3: Coefficients of the $u_{\text{lin}}$ confidence function (model: ember-v1, metric: mAP). Recall that documents are sorted in increasing order of relevance scores. Therefore, document 10 corresponds to the document with highest relevance score while document 1 is the least relevant to the query.
Table 4: nAUCs (in %) for all reference-test relevant pairs (model: ember-v1, metric: mAP).
@@
A simple observation that can be made from Figure 3 is that the value taken by the coefficients depends very strongly on the number of positive documents per instance (see Table A). Indeed, when we look at the results for the SciDocs dataset (five positive documents per instance on average), we see that $u_{\text{lin}}$ has positive weights on the first five documents and negative weights on the last five. A similar observation can be made for the other five datasets.
@@ Appendix D Additional Reference-Based Confidence Functions
In this section, we introduce other functions and approaches to tackle confidence assessment in the reference-based setting.
D.1 Additional Regression-Based Confidence Functions
We propose two additional regression-based confidence functions (Section 3.2): $u_{\text{rf}}$ and $u_{\text{mlp}}$. $u_{\text{rf}}$ is based on a random forest (Ho, 1995) fitted using 100 independent estimators and squared error impurity criterion. 100 was chosen among several values (10, 50, 100, 200, 500, 1000) as the best compromise between efficiency and effectiveness regarding nAUC. $u_{\text{mlp}}$ is based on a Multi-Layer Perceptron (MLP) (Rumelhart et al., 1986) with one hidden layer of size 128 (retained among several values: 32, 64, 128, 256), ReLU activation, and mean squared error loss function. 0.05 was chosen among multiple learning rate values (0.001, 0.005, 0.01, 0.05, 0.1), a batch size equal to that of the reference set was selected (one single iteration per epoch), and 500 training iterations were performed, as it showed the best efficiency-effectiveness trade-off in terms of downstream nAUC.
Results using these two functions are reported in Table 5. We observe that although these two new methods logically outperform the reference-free baselines, they are still less effective than simple linear regression. This is not entirely surprising, given the limited data and constrained feature space at stake in this setup.
D.2 Other Approaches
D.2.1 Classification-Based Approach
@@
$\begin{split}\mathcal{S}_{\bm{\theta}}^{\text{clf}}=\bigr{\{}\left(\mathbf{z},\mathbbm{1}_{\left\{y>m^{+}\right\}}-\mathbbm{1}_{\left\{y<m^{-}\right\}}\right)\ |\ (\mathbf{z},y)\in\mathcal{S}_{\bm{\theta}}\bigl{\}}.\end{split}$ (9)
where $m^{+}>m^{-}\in\mathbb{R}$ are the thresholds above and below which an instance is considered good and bad respectively. Then, we fit a probabilistic classifier $\pi:\mathbb{R}^{k}\rightarrow[0,1]^{3}$ on $\mathcal{S}_{\bm{\theta}}^{\text{clf}}$, taking a vector of sorted scores $s(\mathbf{z})\in\mathbb{R}^{k}$ as input and returning the estimated probabilities of $s(\mathbf{z})$ to belong to classes $-1$ (bad instances), 00 (average instances) and $1$ (good instances). We finally define confidence function $u_{\text{clf}}$ such that for a given unsorted vector of relevance scores $\mathbf{z}$, $u_{\text{clf}}(\mathbf{z})=\pi_{2}\left(s(\mathbf{z})\right)-\pi_{0}\left(s(\mathbf{z})\right)$ (probability that $\mathbf{z}$ stems from a good instance minus the probability that $\mathbf{z}$ stems from a bad instance). The higher $u_{\text{clf}}(\mathbf{z})$, the more confident we are that $\mathbf{z}\in\mathbb{R}^{k}$ stems from a good instance.
In this work, we use a logistic regressor (Cox & Snell, 1989) with $l_{2}$ regularization parameter $\lambda=0.1$, $u_{\text{log}}$, and consider by default that the top and bottom 25% of instances with respect to metric function $m$ represent the good and bad instances respectively. We challenge this parameter in Appendix D.3.2.
D.2.2 Distance-Based Approach
@@
For this purpose, based on thresholds $m^{+}$ and $m^{-}$ defined in Section D.2.1, we build $\mathcal{Z}_{\bm{\theta}}^{+}$ and $\mathcal{Z}_{\bm{\theta}}^{-}$, the sets composed of good and bad instances respectively. Formally,
@@
Then, we construct a confidence scorer $u_{\text{dist}}$ based on the distributions observed in $\mathcal{Z}_{\bm{\theta}}^{+}$ and $\mathcal{Z}_{\bm{\theta}}^{-}$. We define $u_{\text{dist}}=\delta\left(s(\cdot),\mathcal{Z}_{\bm{\theta}}^{-}\right)-\delta\left(s(\cdot),\mathcal{Z}_{\bm{\theta}}^{+}\right)$ as the difference between the distance to the set of bad instances and the distance to the set of good instances. [^12] As for the other reference-based $u$ functions, $u_{\text{dist}}$ takes a vector of unsorted scores as input, sorting being encompassed inside the function. Intuitively, the higher $u_{\text{dist}}(\mathbf{z})$, the more confident we are that $\mathbf{z}$ stems from a good instance.
In this experiment, we rely on the Mahalanobis distance (Mahalanobis, 2018; Lee et al., 2018; Ren et al., 2021). We compute the Mahalanobis distance between set $\mathcal{Z}\subset\mathbb{R}^{k}$ and $\mathbf{z}\in\mathbb{R}^{k}$ as
@@
where $\bm{\mu}_{\mathcal{Z}}\in\mathbb{R}^{k}$ is the average vector of reference set $\mathcal{Z}$ and $\Sigma_{\mathcal{Z}}^{-1}\in\mathbb{R}^{k\times k}$ is the inverse covariance matrix.
@@ D.3 Experimental Results
D.3.1 Abstention Performance
Table 5: nAUCs (in %) averaged over the whole set of available models, for every method including $u_{\text{log}}$ and $u_{\text{mah}}$.
Method $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{mlp}}$ $u_{\text{rf}}$ $u_{\text{log}}$ $u_{\text{mah}}$
@@ D.3.1 Abstention Performance
SciDocs mAP 49.2 58.7 4.7 65.4 65.6 67.3 68.8 65.8
mNDCG 54.9 62.9 10.2 66.6 65.3 68.6 70.2 66.8
mRR 80.2 80.3 41.3 80.5 76.8 70.7 78.3 65.6
AskUbuntu mAP 24.1 16.2 7.3 22.0 19.4 14.1 19.9 10.1
mNDCG 28.1 15.7 8.6 24.5 21.2 14.0 22.0 10.5
mRR 32.7 16.6 14.0 26.2 22.8 14.8 24.1 11.3
StackOverflow mAP 20.7 24.6 35.7 42.6 39.5 37.1 42.0 37.5
mNDCG 21.0 25.0 35.6 42.6 36.0 37.1 42.1 37.4
mRR 21.4 24.8 36.1 42.8 39.6 37.3 42.2 37.7
Alloprof mAP 16.6 17.0 19.3 29.7 25.6 24.3 29.3 25.1
mNDCG 16.8 16.9 18.8 29.3 23.7 24.1 28.9 24.9
mRR 16.9 16.7 19.1 28.9 24.9 23.5 28.6 24.3
CMedQAv1 mAP 24.4 23.3 22.3 30.6 23.3 24.8 29.0 25.7
mNDCG 25.0 23.5 22.3 30.7 22.2 25.4 27.0 24.8
mRR 26.5 24.2 23.4 31.5 24.7 25.5 30.2 25.5
Mmarco mAP 30.1 31.0 39.4 34.0 32.0 29.7 25.1 22.7
mNDCG 30.5 30.9 38.8 34.1 30.3 28.7 25.1 23.3
mRR 30.1 31.0 39.6 34.3 31.8 29.4 23.7 22.8
Average mAP 27.5 28.5 21.5 37.4 34.2 32.9 35.7 31.1
mNDCG 29.4 29.1 22.4 38.0 33.1 33.0 35.9 31.3
mRR 34.6 32.3 28.9 40.7 36.8 33.5 37.9 31.2
@@
D.3.2 Impact of Instance Qualification
Figure 4: nAUC vs. instance quantile threshold for all datasets (model: ember-v1, metric: mAP).
In this section, we analyze the impact of the instance qualification threshold on the performance of $u_{\text{log}}$ and $u_{\text{mah}}$ confidence functions. More precisely, the idea of this experiment is to vary the qualification threshold [^13] As a reminder, a threshold of 10% means that good instances represent the top-10% with regard to the metric of interest, while bad instances represent the bottom-10%. and to test its impact on nAUC over the whole benchmark. For the sake of readability, we present the results for the ember-v1 model and for the mAP only (see Figure 4).